Can you please help me answer this question.
5/y-3 = y+7/2y-6 + 1

Answer by stanbon(75887)

5/y-3 = y+7/2y-6 + 1
---
Multiply thru by 2(y-3) to get:
10 = (y+7) + 2(y-3)
10 = y+7 + 2y - 6
3y +1 = 10
3y = 9
y = 3
---------------
Cheers,
Stan H.